# Beamformed 360° Sound Maps: U‑Net‑Driven Acoustic Source Segmentation and Localization

### v1

@@ Abstract
We introduce a U-net model for 360° acoustic source localization formulated as a spherical semantic segmentation task. Rather than regressing discrete direction-of-arrival (DoA) angles, our model segments beamformed audio maps (azimuth × elevation) into regions of active sound presence. Using delay-and-sum (DAS) beamforming on a custom 24-microphone array, we generate signals aligned with drone GPS telemetry to create binary supervision masks. A modified U-Net, trained on frequency-domain representations of these maps, learns to identify spatially distributed source regions while addressing class imbalance via the Tversky loss. Because the network operates on beamformed energy maps, the approach is inherently array-independent and can adapt to different microphone configurations without retraining from scratch. The segmentation outputs are post-processed by computing centroids over activated regions, enabling robust DoA estimates. Our dataset includes real-world open-field recordings of a DJI Air 3 drone, synchronized with 360° video and flight logs across multiple dates and locations. Experimental results show that U-net generalizes across environments, providing improved angular precision, offering a new paradigm for dense spatial audio understanding beyond traditional Sound Source Localization (SSL).
Index Terms:
sound-source localization, beamforming, semantic segmentation, U-Net, drone acoustics.1 Introduction
@@
SSL is a fundamental task in spatial audio analysis, with applications ranging from surveillance, security, search and rescue, environmental monitoring, and wildlife tracking [1]. SSL is often reduced to estimating the DoA of sound sources. The DoA is usually defined as the azimuth and elevation angle of the direction of the audio source while ignoring the distance to it. Microphone arrays can be steered to act as spatial filters, which enables manipulation of the array’s directivity (also referred to as beamforming).[2]. Traditional SSL techniques rely on signal processing algorithms such as time-difference of arrival (TDOA), generalized cross-correlation with phase transform (GCC-PHAT)[3], Steered-response power-phase transform (SRP-PHAT)[4] and multiple signal classification (MUSIC)[5], or beamforming methods like (DAS)[6]. While effective in controlled environments, these approaches degrade in performance under noise, reverberation, or when dealing with moving sources and complex acoustic scenes. Recent deep learning methods have achieved notable improvements in accuracy and robustness, even in challenging scenarios with noise, reverberation, and multiple simultaneous sources [7]. In recent years, the use of multichannel audio processing and visual perception has gained traction [8], particularly with the use of microphone arrays and beamforming to create spatial energy maps [9] essentially turning sound into images, some of those systems are known as acoustic cameras.[10].
@@
In this work, we propose a novel spherical segmentation framework for localizing sound sources using microphone arrays. Inspired by the image recognition paradigm in computer vision like YOLO [11] or Deeplapv3[12], we develop a U-Net-based architecture[13] that performs binary segmentation over a 2D spherical acoustic image (azimuth × elevation) derived from DAS beamforming. Instead of regressing point estimates of direction, our model learns to segment the acoustic field, highlighting regions in space where sound sources are present. This formulation enables spatial mapping of targets such as drones. To enable supervised learning, we construct and release a labeled dataset based on real-world recordings of a DJI Air 3 drone in open-field conditions on different days and locations. The public dataset comprises 24‑channel audio, GPS‑aligned binary masks and 360° video.
@@
The main contributions of this paper include the introduction of a novel paradigm for acoustic source localization formulated as a spherical acoustic image segmentation task, inspired by semantic segmentation approaches in computer vision. The authors design and train a U-Net-based model, which performs segmentation of sound source directions on beamformed audio maps. To support further research in audio localization and learning-based sound source localization (SSL), they also release a unique multi-channel, real-world dataset that includes synchronized audio, spatial labels, and drone telemetry.
@@ 2 Related Work
Acoustic perception tasks explicitly distinguish three sequential objectives: detection, classification, and localization. Many classical SSL systems tackle them in isolation rather than in an integrated pipeline [1]. Recent Sound Event Detection and Localisation (SELD) systems, also referred to as SSL, encompass two tasks: sound event detection (SED) and DoA, which are separate outputs of the neural network. The SED branch performs a multi-label classification task, and the DoA branch performs a multi-output regression task, as cited in [14] and [15].
Our work resolves SELD as a semantic-segmentation problem; the resulting mask simultaneously answers: (i) is there a source? (detection), (ii) where is it? (localisation) and through class specific training (iii) what type is it? (classification).
@@ 3.4 Beamforming Map Construction and Spectral Feature Extraction
In general, beamforming relies on the far-field assumption that the sources are far away and the waves become spherical or planar at the array position [2]. Given the position vector of each microphone in the array, the delay is obtained via the projection of onto the direction vector:
@@ 3.7 U-Net Architecture and Hyperparameter Optimization
To address label sparsity and class imbalance, we selected the Tversky loss [32] which extends Dice loss by weighting false positives and is effective for small positive regions.
@@ 5 Conclusion and Future Work
This work presents a U-net framework for sound source localization (SSL) that reinterprets beamformed acoustic maps as spatial segmentation tasks over the spherical field. By applying U-Net-based convolutional architectures to azimuth-elevation representations of DAS beamformed audio, we address classical limitations in beamforming—such as low-frequency blurring and high-frequency side lobes—thus enhancing angular resolution and robustness. Transforming beamformed maps into polar coordinates further aligns the spatial layout with spherical geometry, reducing distortion and better supporting CNN-based learning, as emphasized in DeepWave [10].
Unlike end-to-end models requiring raw microphone inputs or fixed array geometries, U-net model applies preprocessed spatial inputs, enabling different mic-array configurations. Combined with real-world drone recordings and GPS-based supervision, the proposed method demonstrates generalization across distances and locations, offering a practical solution for low-latency acoustic perception.
Future work will focus on extending the system to handle spatio-temporal dynamics via recurrent or attention-based models like Mark R-CNN[33] or yolact++[34] and VisTR[35], supporting multi-source and multiclass segmentation, and using synthetic data generation pipelines for diverse acoustic environments.
By framing SSL as a semantic segmentation task on beamformed maps, this approach bridges spatial signal processing and computer vision, opening new directions for high-resolution, real-time acoustic scene understanding.

### v2

# Acoustic Imaging for Low-SNR UAV Detection: Dense Beamformed Energy Maps and U-Net SELD

@@ Abstract
We introduce a U-net model for 360° acoustic source localization formulated as a spherical semantic segmentation task. Rather than regressing discrete direction-of-arrival (DoA) angles, our model segments beamformed audio maps (azimuth × elevation) into regions of active sound presence. Using delay-and-sum (DAS) beamforming on a custom 24-microphone array, we generate signals aligned with drone GPS telemetry to create binary supervision masks. A modified U-Net, trained on frequency-domain representations of these maps, learns to identify spatially distributed source regions while addressing class imbalance via the Tversky loss. Because the network operates on beamformed energy maps, the approach is inherently array-independent and can adapt to different microphone configurations and can be transferred to different microphone configurations with minimal adaptation. The segmentation outputs are post-processed by computing centroids over activated regions, enabling robust DoA estimates. Our dataset includes real-world open-field recordings of a DJI Air 3 drone, synchronized with 360° video and flight logs across multiple dates and locations. Experimental results show that U-net generalizes across environments, providing improved angular precision, offering a new paradigm for dense spatial audio understanding beyond traditional Sound Source Localization (SSL). We additionally validate the same beamforming-plus-segmentation formulation on the DCASE 2019 TAU Spatial Sound Events benchmark, showing that the approach generalizes beyond drone acoustics to multiclass Sound Event Localization and Detection (SELD) scenarios.
1 Introduction
@@
SSL is a fundamental task in spatial audio analysis, with applications ranging from surveillance, security, search and rescue, environmental monitoring, and wildlife tracking [19]. SSL is often reduced to estimating the DoA of sound sources. The DoA is usually defined as the azimuth and elevation angle of the direction of the audio source while ignoring the distance to it. Microphone arrays can be steered to act as spatial filters, which enables manipulation of the array’s directivity (also referred to as beamforming).[4]. Traditional SSL techniques rely on signal processing algorithms such as time-difference of arrival (TDOA), generalized cross-correlation with phase transform (GCC-PHAT)[8], Steered-response power-phase transform (SRP-PHAT)[12] and multiple signal classification (MUSIC)[27], or beamforming methods like (DAS)[21]. While effective in controlled environments, these approaches degrade in performance under noise, reverberation, or when dealing with moving sources and complex acoustic scenes. Recent deep learning methods have achieved notable improvements in accuracy and robustness, even in challenging scenarios with noise, reverberation, and multiple simultaneous sources [13]. In outdoor long-range scenarios, localization performance is frequently SNR-limited by propagation and non-stationary ambient noise. As distance increases, high-frequency UAV components attenuate rapidly and background noise becomes partially spatially coherent (e.g., wind, vegetation, traffic), reducing effective array gain. This motivates hybrid designs where a computationally simple beamformer produces an interpretable spatial map, and learning is used to denoise and regularize the map structure rather than operating directly on raw multichannel waveforms. In recent years, the use of multichannel audio processing and visual perception has gained traction [14], particularly with the use of microphone arrays and beamforming to create spatial energy maps [31] essentially turning sound into images, some of those systems are known as acoustic cameras.[28].
@@
In this work, we propose a novel spherical segmentation framework for localizing sound sources using microphone arrays. Inspired by the image recognition paradigm in computer vision like YOLO [23] or Deeplapv3[9], we develop a U-Net-based architecture[24] that performs binary segmentation over a 2D spherical acoustic image (azimuth × elevation) derived from DAS beamforming. Instead of regressing point estimates of direction, our model learns to segment the acoustic field, highlighting regions in space where sound sources are present. This formulation enables spatial mapping of targets such as drones. To enable supervised learning, we construct a labeled dataset based on real-world recordings of a DJI Air 3 drone in open-field conditions on different days and locations. The dataset will be made publicly available upon journal publication and includes 24‑channel audio, GPS‑aligned binary masks and 360° video.
@@
The main contributions are:
Segmentation formulation for SSL/SELD: we re-parameterize localization as dense spatial mask prediction over azimuth–elevation maps, enabling joint detection and localization via a single output representation.
Beamforming-to-image pipeline for 360∘ maps: we build frequency-binned beamformed map tensors and introduce a polar reprojection for hemisphere coverage to better match convolutional spherical distortions.
Real open-field UAV dataset with GPS supervision: we provide a field dataset with synchronized multichannel audio and drone telemetry for learning spatial segmentation under realistic outdoor recordings.
Cross-domain validation on TAU Spatial: we re-cast DCASE 2019 Task 3 into multiclass spatial segmentation and demonstrate that the same formulation generalizes beyond drones.
@@ 2 Related Work
Acoustic perception tasks explicitly distinguish three sequential objectives: detection, classification, and localization. Many classical SSL systems tackle them in isolation rather than in an integrated pipeline [19]. Recent Sound Event Detection, Localisation and Classification(SELD/SELC) systems, also referred to as SSL, encompass two tasks: sound event detection (SED) and DoA, which are separate outputs of the neural network. The SED branch performs a multi-label classification task, and the DoA branch performs a multi-output regression task, as cited in [26] and [1].
Our work resolves SELD/SELC as a semantic-segmentation problem; the resulting mask simultaneously answers: (i) is there a source? (detection), (ii) where is it? (localisation) and through class specific training (iii) what type is it? (classification).
Conceptually, this paper keeps the SELD/SE objective but changes the output parameterization: instead of predicting class-wise event activity and DoA regression, we predict class-wise spatial occupancy masks over an azimuth–elevation grid. DoA estimates are then obtained by simple geometric post-processing (centroiding).
@@ 2.3 Gap Analysis and Motivation
- A multiclass validation on DCASE 2019 TAU Spatial Sound Events, showing that the segmentation formulation remains competitive in DoA accuracy beyond drone data.
@@ 3.4 Beamforming Map Construction and Spectral Feature Extraction
In general, beamforming relies on the far-field assumption that the sources are far away and the waves become planar at the array position [4]. Given the position vector of each microphone in the array, the delay is obtained via the projection of onto the direction vector:
@@ 3.7 U-Net Architecture and Hyperparameter Optimization
We use task-specific loss functions depending on label sparsity and class structure. For binary drone segmentation we employ the Tversky loss to explicitly control the FP/FN trade-off under severe foreground sparsity. For multiclass TAU Spatial we use a multilabel Dice + BCE-with-logits objective with per-class positive weights computed from training pixels, which stabilizes learning when class frequencies differ markedly.
@@ 4.1 Performance Analysis by Distance Range
Comparison against classical SSL baselines (Test 1). Figure 6 compares the proposed U-Net against representative classical SSL pipelines on the unseen Test 1 condition.
For each method, a position estimate is extracted per frame and compared to the GPS-derived ground truth using the spherical angular distance, results are aggregated by distance range.
Across short-to-mid ranges, the classical baselines exhibit comparable trends, but their angular error increases at longer distances (approximately 90 m), consistent with reduced effective SNR and increasing sensitivity to interference and sidelobe structure in the beamformed maps. In contrast, the U-Net maintains substantially lower error over the same distance, indicating that is able to suppress spurious peaks. While the U-Net error also increases with distance, its degradation is markedly slower, supporting the conclusion that learning a spatial mask over the full azimuth-elevation field provides improved robustness compared to selecting a single maximum-response direction.
4.2 Multiclass Evaluation on the DCASE 2019 TAU Spatial Sound Events Dataset
To further validate the generality of the proposed segmentation-based formulation, we evaluated the model on the DCASE 2019 Task 3 TAU Spatial Sound Events dataset (Task 3)[2], a widely adopted benchmark for multiclass (SELD).
4.2.1 Dataset and Preprocessing
The TAU Spatial dataset consists of synthetically generated spatial sound scenes to a tetrahedral four-microphone array. Each recording contains between one and two simultaneously active sound events, positioned at a fixed distance of 1 m and 2 m from the array, with a SNR of approximately 30 dB. The dataset provides time-aligned annotations in azimuth and elevation for multiple sound event classes.
To ensure methodological consistency with the proposed drone pipeline, the dataset was reprocessed into beamformed acoustic maps and converted into a multiclass semantic segmentation task. A regular tetrahedral array geometry was assumed, with an effective radius of 4.2 cm. For each 100 ms audio chunk, delay-and-sum (DAS) beamforming was applied over a discrete azimuth–elevation grid. This produced a three-dimensional beamformed tensor indexed by azimuth, elevation, and time.
Each beamformed time-domain signal was transformed into the frequency domain using the real-valued FFT. We preserve strict parity with the feature extraction pipeline used in the drone experiments. Magnitude spectra were extracted, restricted to the frequency range 100–10,000 Hz, and globally normalized using z-score statistics computed over the training set. The normalized spectra were then grouped into contiguous frequency bins (we use for TAU Spatial and for UAV). In the TAU Spatial case, we do not apply the polar reprojection used in the UAV pipeline: the elevation annotations are confined to a limited range ( to ), so the azimuth–elevation mask remains well-conditioned on the rectangular grid without the severe geometric warping observed in hemispherical maps. In contrast, the UAV experiments span the upper hemisphere ( to elevation), where sampling and label geometry become increasingly distorted near the zenith, motivating the polar (radial-elevation) mapping adopted for the drone setting.
To ensure methodological consistency with the proposed drone pipeline, the dataset was reprocessed into DAS beamformed acoustic maps and converted into a multiclass semantic segmentation task, as summarized in Fig. 7.
4.2.2 Multiclass Label Construction
For each time frame, class-specific binary segmentation masks were generated by comparing the ground-truth event direction with every steering direction in the azimuth–elevation grid. Grid points within an angular tolerance of from the annotated direction were labeled as active for the corresponding class. This procedure resulted in one binary mask per sound class.
4.2.3 Network Architecture and Training Protocol
The same U-Net architecture used in the UAV experiments was employed for the DCASE benchmark, with minor adaptations for multiclass output. The network processes beamformed acoustic maps and predicts a dense segmentation mask with one output channel per sound class. Training was performed in a multilabel setting using a combined Dice and binary cross-entropy loss. To compensate for severe class imbalance and sparse spatial labels, per-class positive weights were computed from the training set based on the ratio of negative to positive pixels. Model selection and early stopping were driven by the macro-averaged F1 score across classes.
4.2.4 Results and Discussion
The segmentation-based U-Net achieves competitive performance on the TAU Spatial dataset despite operating on beamformed energy maps rather than raw multichannel signals.
Fig. 8 shows a qualitative example of the proposed map-to-mask multiclass pipeline on TAU Spatial. Starting from a single-frame DAS beamformed map, the U-Net outputs class-specific regions in the azimuth–elevation grid, effectively transforming a broad and often ambiguous beamformer response into a structured semantic mask. Occasional multiple disconnected detections within the same class can arise due to the limited aperture of the tetrahedral four-microphone array and the sidelobe structure of DAS, which can yield competing spatial hypotheses within a single frame. This observation motivates temporal context (multi-frame input) and/or lightweight spatial association to improve robustness under overlapping events and transient classes. Table 1 situates the proposed pipeline against established SELD baselines on TAU Spatial 2019. On the full evaluation condition (up to two overlapping sources), the segmentation-based U-Net presents an overall F-score of with a mean DoA error of , improving to F-score and DoA error when restricting to frames containing a single active source. In contrast, strong end-to-end SELD systems such as SELDnet and w2v-SELD report 94–95% F-score with 3.7–4.7∘ DoA error, while the DCASE baseline SELDnet exhibits a pronounced degradation in localization error (38.1∘). Overall, these results support that reframing SELD as dense semantic segmentation over spatial acoustic maps can preserve competitive localization accuracy, while the drop under overlapping sources is consistent with the lack of explicit temporal association and multi-instance tracking in a purely frame-wise segmentation formulation. To address this limitation, we are currently developing a temporal variant that ingests multiple consecutive map frames (i.e., increased temporal receptive field) to better capture event dynamics and improve F-score, particularly on TAU Spatial where many classes are impulsive/transient (e.g., knock, drawer, doorslam, keyboard) and thus benefit from temporal context. This differs from the original UAV-driven design target, where the acoustic signature is comparatively sustained and temporally stable, making single-frame segmentation more viable. Finally, we note that the ground-truth mask design should be consistent with the effective array resolution: with fewer microphones the beamformer main lobe is wider (lower angular resolution), which motivates using a larger angular mask tolerance, whereas denser/larger-aperture arrays can support a tighter (smaller) mask due to a narrower main lobe.
| Model | F1 (%) | DoA Error (∘) |
| SELDnet [16] | 94.7 | 3.7 |
| Baseline SELDnet [2] | 83.1 | 38.1 |
| w2v-SELD [26] | 94.7 | 4.7 |
| U-Net (eval, up to two sources) | 68.0 | 4.9 |
| U-Net (one-source subset) | 79.0 | 3.0 |
These results demonstrate that reframing SELD as a dense semantic segmentation problem over spatial acoustic maps is not limited to drone acoustics, but generalizes to established multiclass benchmarks. Overall, this experiment confirms that the proposed beamforming-plus-segmentation paradigm provides a possible flexible and array-agnostic alternative to conventional SELD architectures, while maintaining competitive localization accuracy and robust detection behavior across domains.
@@ 5 Conclusion and Future Work
This work presents a U-Net framework for sound source localization (SSL) that reinterprets beamformed acoustic maps as spatial segmentation tasks over the spherical field. By applying U-Net-based convolutional architectures to azimuth–elevation representations of DAS beamformed audio, we address classical limitations in beamforming—such as low-frequency blurring and high-frequency side lobes—thus enhancing angular resolution and robustness. Transforming beamformed maps into polar coordinates further aligns the spatial layout with spherical geometry, reducing distortion and better supporting CNN-based learning, as emphasized in DeepWave [28].
Unlike end-to-end models requiring raw microphone inputs or fixed array geometries, the proposed U-Net model operates on preprocessed spatial representations, enabling transfer across different microphone array configurations. Combined with real-world drone recordings and GPS-based supervision, the method demonstrates robust generalization across distances and recording conditions, providing a practical solution for low-latency outdoor acoustic perception.
Beyond the UAV scenario, we further validated the proposed beamforming-plus-segmentation paradigm on the DCASE 2019 TAU Spatial Sound Events dataset, a standard benchmark for multiclass SELD. Despite relying on beamformed energy maps rather than raw multichannel waveforms, the segmentation-based formulation achieves competitive direction-of-arrival accuracy when compared to established SELDnet-style architectures. This result confirms that dense spatial mask prediction constitutes a viable alternative parameterization of SELD, in which detection and localization emerge jointly from semantic segmentation of the acoustic field.
Future work will focus on extending the framework to handle spatio-temporal dynamics via recurrent or attention-based models such as Mask R-CNN [35], YOLACT++ [6], or VisTR [32], enabling explicit temporal association, multi-source instance tracking, and improved performance under overlapping sound events. Additional directions include synthetic data generation for broader acoustic diversity and multi-array fusion for scalable real-time deployments.
By framing SSL and SELD as semantic segmentation tasks on beamformed spatial representations, this work bridges classical array signal processing and modern computer vision, opening new directions for high-resolution, interpretable, and deployable acoustic scene understanding.